OpenStudy (anonymous):

What Is the Solution Of The System? Use Elimination: x=y-6 6x-37=-9 Please Help Me, I Don't Quite Get It. . . =)

OpenStudy (anonymous):

6x - 37 = -9 6x = 28 x = 14/3 x + 6 =y y = 32/3

OpenStudy (anonymous):

Thank You So Much!